# Interferometry of non-Abelian band singularities and Euler class topology

### v1

@@ Acknowledgements.
where $c\equiv\cos{\phi_{I}},s\equiv\sin{\phi_{I}}$ and $\phi_{I}$ is the angle by which the frame rotates between ingoing and outgoing paths (see Fig. 1(c)). The phase $\alpha_{n}^{I}$ consists of dynamic and geometric contributions (which are changes by either 00 or $\pi$) and from the pre- (i) and post-node (ii) sections of the path, $\alpha_{n}^{I}=\alpha_{n}^{I}(i)+\alpha_{n}^{I}(ii)$, as demonstrated in Fig. S5. $\alpha_{n+1}^{I}$ consists of the phase developed in band $n$ before the node, and the dynamic/geometric phases accumulated in band $n+1$ after the node, $\alpha_{n+1}^{I}=\alpha_{n}^{I}(i)+\alpha_{n+1}^{I}(ii)$. Along path II we similarly have the final wavefunction
@@
Here, the relative directions of frame winding $\phi_{II}=\pm\phi_{I}$ leads to a relative sign in the component excited at the node. Different than the main text, we can here assume that this frame winding is encompassed in the final phase $\alpha^{II}_{n+1}$ for compactness.
@@
In the wavefunctions above, we have assumed that the magnitude of the frame rotation angle $\phi$ is the same at both nodes, $|\phi_{I}|=|\phi_{II}|$. By studying the winding about a single node using e.g. the techniques of Ref. [34], the turning angles $\theta$ at each node can be chosen such that this is satisfied, making our interferometry generically applicable even under reduced lattice symmetries.
@@
.2.1 Accounting for Dynamic Phases
Whilst the phases above consist of both a dynamic part $\alpha^{\prime}$ and a geometric part $\alpha=\alpha^{geo}$, $\Delta\alpha_{n}^{\mathrm{tot}}=\Delta\alpha_{n}+\Delta\alpha_{n}^{\prime}$, our goal here is to extract the geometric part only. Here, we discuss how this can be achieved. Upon performing the interferometry, the final spin-up population is [Eqs. (S6,S7)]
$p_{\uparrow}=\cos^{2}{(\phi)}\cos^{2}{\left(\Delta\alpha_{n}^{\mathrm{tot}}/2\right)}+\sin^{2}{(\phi)}\cos^{2}{\left(\Delta\alpha_{n+1}^{\mathrm{tot}}/2\right)}.$ (S10)
By performing this interferometry at two (known) values of $\phi$, or by measuring spin populations separately in bands $n$ and $n+1$ [45], this allows for $\cos^{2}{\left(\Delta\alpha_{n}^{\mathrm{tot}}/2\right)}$ and $\cos^{2}{\left(\Delta\alpha_{n+1}^{\mathrm{tot}}/2\right)}$ to be determined. We now seek to determine the geometric contribution, which may be either 00 or $\pi$, to these phases.
We here construct a ‘reference interferometry loop’ technique which relies on repeating the interferometry with twice the acceleration, halving the time for the total process. As a result, the dynamic phase contribution is halved, whilst the geometric phase remains fixed, therefore determining $\cos^{2}{\left(\Delta\alpha^{\prime}_{n(+1)}/4+\Delta\alpha_{n(+1)}/2\right)}$.
The problem is then reduced to determining whether $\alpha=0$ or $\pi$ when we have $x=\cos^{2}\left(\alpha^{\prime}/2+\alpha/2\right)$ and $y=\cos^{2}\left(\alpha^{\prime}/4+\alpha/2\right)$ (we here drop the $\Delta$ and subscripts for clarity). This can be achieved e.g. by calculating $x-(2y^{2}-1)^{2}=\sin{\left(\alpha/2\right)}\sin{\left(\alpha^{\prime}+3\alpha/2\right)}$. Therefore when $\alpha=0$, $x-(2y^{2}-1)^{2}=0$, and generally this quantity will be non-zero when $\alpha=\pi$. We note that when $\alpha^{\prime}=\pm\pi/2$, this will yield an ‘accidental zero’. This can be resolved by performing the reference interferometry with four times the acceleration, yielding $z=\cos^{2}\left(\alpha^{\prime}/8+\alpha/2\right)$, and $y-(2z^{2}-1)^{2}=\sin{\left(\alpha/2\right)}\sin{\left(\alpha^{\prime}/2+3\alpha/2\right)}$ can be calculated. If this is also 00, then one can be certain that the geometric phase $\alpha=0$; if we previously had an ‘accidental zero’, then we expect $1/\sqrt{2}$. Indeed, this set of measurements provides a reasonably robust answer. This is because $\mathrm{min}(\mathrm{abs}(x-(2y^{2}-1)^{2})+\mathrm{abs}(y-(2z^{2}-1)^{2}))=1/\sqrt{2}$ when $\alpha=\pi$, providing a significant ‘gap’ to the value expected for $\alpha=0$ (i.e. 00).
@@
.3 III. Details for Method 2: Consecutive Deflection
Suppose that we start in a state purely in band $\mathinner{|{u_{n}}\rangle}$. When we evolve this in and out of a node at an angle such that the frame has rotated by $\phi_{1}$, upon exiting the node the state is given by
$\cos{\phi_{1}}\mathinner{|{u_{n}}\rangle}-\sin{\phi_{1}}\mathinner{|{u_{n+1}}\rangle}.$ (S11)
@@
$e^{i\alpha^{\prime}}\cos{\phi_{1}}\mathinner{|{u_{n}}\rangle}-\sin{\phi_{1}}\mathinner{|{u_{n+1}}\rangle}.$ (S12)
Upon entering and exiting this second node with a frame rotation angle $\phi_{2}$, we have (denoting $\cos{\phi_{i}}=c_{i},\sin{\phi_{i}}=s_{i}$)
@@
$n_{1}=|e^{i\alpha}c_{1}c_{2}-s_{1}s_{2}|^{2}=\cos^{2}{\left(\phi_{1}+\phi_{2}\right)}+\sin{2\phi_{1}}\sin{2\phi_{2}}\sin^{2}{\left(\alpha/2\right)}.$ (S14)
@@
$n_{1}$ $=\frac{1}{2}\left[\cos{\left(2\phi_{1}+2\phi_{2}\right)}+1+2\sin{2\phi_{1}}\sin{2\phi_{2}}\sin^{2}{\left(\alpha/2\right)}\right]$ (S15)
$=\frac{1}{2}\left[\cos{2\phi_{1}}\cos{2\phi_{2}}-\cos{\alpha}\sin{2\phi_{1}}\sin{2\phi_{2}}+1\right]$ (S16)
$=\frac{1}{2}\left[A\cos{\left(2\phi_{1}-\beta\right)}+1\right]$ (S17)
@@
$A^{2}=\cos^{2}{\left(2\phi_{2}\right)}+\cos^{2}{\left(\alpha\right)}\sin^{2}{\left(2\phi_{2}\right)},$ (S18)
$\tan{\beta}=-\cos{\left(\alpha\right)}\tan(2\phi_{2})$ (S19)
@@
.4 IV. Numerical calculations
Here we describe the numerical validation for the time-evolution along paths in the BZ.
When an external force $\vec{F}$ is applied to particle which otherwise has Hamiltonian $H$, the Schrödinger equation is
$i\partial_{t}\mathinner{|{\psi}\rangle}=(H-\vec{F}\cdot\vec{r})\mathinner{|{\psi}\rangle}.$ (S20)
$\mathinner{|{\psi}\rangle}=e^{i\bm{k}\cdot\vec{r}}\mathinner{|{u_{k}}\rangle},$ (S21)
where $\mathinner{|{u_{k}}\rangle}$ is a superposition of bands
$\mathinner{|{u_{k}}\rangle}=\sum_{n}\alpha_{n}\mathinner{|{u_{n,k}}\rangle}.$ (S22)
The effect of the force is that $\dot{\bm{k}}=\vec{F}$ [50], such that through an applied force we can choose the path through momentum-space $\bm{k}(t)$. Therefore, the wavefunction will evolve as
$\mathinner{|{\psi(t)}\rangle}=e^{i\bm{k}(t)\cdot\vec{r}}\mathinner{|{u_{k(t)}}\rangle},$ (S23)
with
$\mathinner{|{u_{k(t)}}\rangle}=\sum_{n}\alpha_{n}(t)\mathinner{|{u_{n,k(t)}}\rangle}.$ (S24)
$i\partial_{t}\mathinner{|{u(t)}\rangle}=H_{k(t)}\mathinner{|{u(t)}\rangle},$ (S25)
where
$H_{k}=e^{-i\bm{k}\cdot\vec{r}}He^{i\bm{k}\cdot\vec{r}}$ (S26)
is the Bloch Hamiltonian, and $\bm{k}(t)$ now simply acts as a varying external parameter.
Defining the path of interest $\bm{k}(t)$, we solve this equation numerically using Runge-Kutta [51], for a given starting state. The initial state is typically a pure band eigenvector.
To confirm the results of the interferometry, we must find the phase difference along two paths for a given band. If the final states along paths $I$ and $II$ are given by $\mathinner{|{u_{I}}\rangle}$ and $\mathinner{|{u_{II}}\rangle}$, and we would like to find the phase difference of the component in band $\mathinner{|{u_{n}}\rangle}$, we evaluate the argument of $\mathinner{\langle{u_{I}|u_{n}}\rangle}\mathinner{\langle{u_{n}|u_{II}}\rangle}$. For example, $\Delta\alpha_{n+1}\Delta\alpha_{n}=\mathinner{\langle{u_{I}|u_{n}}\rangle}\mathinner{\langle{u_{n}|u_{II}}\rangle}-\mathinner{\langle{u_{I}|u_{n+1}}\rangle}\mathinner{\langle{u_{n+1}|u_{II}}\rangle}$. The populations were calculated from these using Eqns. S6-S9. We worked backwards from these populations, using the techniques to deal with dynamic phases that were described in both the main text (for Method 2) and supplementary materials (for Method 1) to recover the geometric phase contributions. For both methods, to achieve adiabaticity we assumed timescales of $t\gtrapprox 10^{2}J^{-1}$, which typically corresponded to a total phase evolution $\mathcal{O}(10^{2})$.
.5 IV A. Results and Examples for Method 1
@@
We performed numerical calculations to verify that the phases and band populations are as expected from the theory. The results, in addition to detailed explanations to clarify the physics, are presented here.
.5.1 Static Kagome with NN and NNN tunneling terms
@@
.5.2 Anomalous Floquet Phases
@@
To deal with the dynamic phases numerically, we employ two approaches. First, we calculate directly the dynamic phase by integrating the band energies along the paths and subtract it at the end to leave only the geometric contribution. We also work directly from the calculated populations, using the reference interferometry technique described in Sec.II ”Accounting for Dynamic Phases” chapter, to extract the geometric phase difference. Both methods are numerically confirmed to allow for determining the correct geometric phase difference.
@@
.6 IV B: Results and Examples for Method 2
In this section we present corroborating numerical results for Method 2. In particular, we take the opportunity to step-by-step demonstrate how our techniques work when the winding is not uniform, and emphasise the distinction between the frame-rotation angle $\phi$ and the path rotation angle $\theta$.
We have confirmed the technique for nearest-neighbour Kagome in Fig. 3 for similarly- and opposite-charged nodes (K and K’). We first note that the value of $\cos{\alpha^{\prime}}$ extracted from the ‘double acceleration’ plots (Fig. 3c) matches up with dynamic phases found by direct integration of energies along the paths, as well as the shifts in the oscillations (Fig. 3b) are quantitatively as expected. In this case, the frames wind uniformly ($\phi=\pm\theta$) and thus choosing $\theta=\pi/2$ fixes $|\phi|=\pi/4$ as required.
@@
Step 1: We first must determine the path angles $\theta$ that lead to frame rotation angles $|\phi|=\pi/4$. To do this, we use the methods of Ref. [34]. Taking the path which joins the nodes as the fixed point, study the population changes upon entering and exiting each node at an angle. This yields the winding shown in Fig. S8. From this data, we can read off the values of $\theta$ for which the population is $0.5$, corresponding to $|\phi|=\pi/4$. In this case, we obtain $\theta_{1}=2.43$ or $\theta_{1}=-0.71$. Repeating this for the other node, we have $\theta_{2}=2.43$ (we do not need the negative option here).
Figure S8: The population in band 1 after entering and exiting the node at an angle $\theta_{1}$ relative to the path joining the nodes, which corresponds to the frame rotation angle $\phi_{1}$ through $p_{1}=\cos^{2}\phi_{1}$. This allows the value of $\theta_{1}$ to be chosen such that the frame rotation angle $|\phi|=\pi/4$. Notice that the winding is not uniform, but the net winding is $\pi$ as it is a topologically protected value.
Step 2: For both these values of $\theta_{1}$, and the fixed $\theta_{2}$, we prepare a wavepacket in band 1 and accelerate it along the required path with an acceleration $2a$ (where $a$ must be chosen such that the evolution is adiabatic), before measuring the final population in band 1. For the two values of $\theta_{1}$ we obtain $p_{1}=0.214$ and $p_{1}=0.786$. The square of the difference of these values gives us the dynamic phase $\cos^{2}{\alpha/2}=0.326$.
Step 3: Finally, we set $\theta_{1}=2.43$ and $\theta_{2}=2.43$ (i.e. the $\theta$’s have the same sign), and accelerate the wavepacket along the path with an acceleration $a$. We find that the final population in band 1 is $0.326$, which is the same as $\cos^{2}{\alpha/2}$, and therefore we know that along this line joining the nodes, the nodes are oppositely charged, and the patch Euler class along this line is 0.

### v2

@@ Acknowledgements.
where $c\equiv\cos{\phi^{\mathrm{frame}}_{I}},s\equiv\sin{\phi^{\mathrm{frame}}_{I}}$ and $\phi^{\mathrm{frame}}_{I}$ is the angle by which the frame rotates between ingoing and outgoing paths (see Fig. 1c). The phase $\alpha_{n}^{I}$ consists of dynamic and geometric contributions (which are changes by either 00 or $\pi$) and from the pre- (i) and post-node (ii) sections of the path, $\alpha_{n}^{I}=\alpha_{n}^{I}(i)+\alpha_{n}^{I}(ii)$, as demonstrated in Fig. S7. $\alpha_{n+1}^{I}$ consists of the phase developed in band $n$ before the node, and the dynamic/geometric phases accumulated in band $n+1$ after the node, $\alpha_{n+1}^{I}=\alpha_{n}^{I}(i)+\alpha_{n+1}^{I}(ii)$. Along path II we similarly have the final wavefunction
@@
Here, the relative directions of frame winding $\phi^{\mathrm{frame}}_{II}=\pm\phi^{\mathrm{frame}}_{I}$ leads to a relative sign in the component excited at the node. Different than the main text, we can here assume that this frame winding is encompassed in the final phase $\alpha^{II}_{n+1}$ for compactness. It is also worth noting here that, in practice, the paths are brought together by applying a $\pi$-pulse at the nodes, which switches the spin indices. However, this is just equivalent to a relabelling of the paths, and so don’t consider this in the main text for simplicity.
@@
In the wavefunctions above, we have assumed that the magnitude of the frame rotation angle $\phi^{\mathrm{frame}}$ is the same at both nodes, $|\phi^{\mathrm{frame}}_{I}|=|\phi^{\mathrm{frame}}_{II}|$. By studying the winding about a single node using e.g. the techniques of Ref. [34], the turning angles $\phi^{\mathrm{frame}}$ at each node can be chosen such that this is satisfied, making our interferometry generically applicable even under reduced lattice symmetries.
@@
.3.1 Accounting for Dynamic Phases
Whilst the phases above consist of both a dynamic part $\alpha^{\prime}$ and a geometric part $\alpha=\alpha^{geo}$, $\Delta\alpha_{n}^{\mathrm{tot}}=\Delta\alpha_{n}+\Delta\alpha_{n}^{\prime}$, our goal is to extract the geometric part only. When there is sufficient symmetry (for example for the paths shown in Fig. 2 of the main text), the dynamic phase evolved is the same along both paths and don’t appear in $\Delta\alpha_{n}^{\mathrm{tot}}$. More generally, however, we must account for their presence. Once potential method is described here.
By measuring spin populations separately in bands $n$ and $n+1$ [46], $\cos^{2}{\left(\Delta\alpha_{n}^{\mathrm{tot}}/2\right)}$ and $\cos^{2}{\left(\Delta\alpha_{n+1}^{\mathrm{tot}}/2\right)}$ can be determined. This can also be achieved by measuring spin populations only, if we separately allow an additional $\pi$ phase shift to develop between the two paths (for example by applying a uniform magnetic field) before the final $\pi/2$-pulse. We now seek to determine the geometric contribution, which may be either 00 or $\pi$, to these phases.
We construct a ‘reference interferometry loop’ technique which relies on repeating the interferometry with twice the acceleration, halving the time for the total process. As a result, the dynamic phase contribution is halved, whilst the geometric phase remains fixed, therefore determining $\cos^{2}{\left(\Delta\alpha^{\prime}_{n(+1)}/4+\Delta\alpha_{n(+1)}/2\right)}$.
The problem is then reduced to determining whether $\alpha=0$ or $\pi$ when we have $x=\cos^{2}\left(\alpha^{\prime}/2+\alpha/2\right)$ and $y=\cos^{2}\left(\alpha^{\prime}/4+\alpha/2\right)$ (we here drop the $\Delta$ and subscripts for clarity). This can be achieved e.g. by calculating $x-(2y-1)^{2}=\sin{\left(\alpha/2\right)}\sin{\left(\alpha^{\prime}+3\alpha/2\right)}$. Therefore when $\alpha=0$, $x-(2y-1)^{2}=0$, and generally this quantity will be non-zero when $\alpha=\pi$. We note that when $\alpha^{\prime}=\pm\pi/2$, this will yield an ‘accidental zero’. This can be resolved by performing the reference interferometry with four times the acceleration, yielding $z=\cos^{2}\left(\alpha^{\prime}/8+\alpha/2\right)$, and $y-(2z-1)^{2}=\sin{\left(\alpha/2\right)}\sin{\left(\alpha^{\prime}/2+3\alpha/2\right)}$ can be calculated. If this is also 00, then one can be certain that the geometric phase $\alpha=0$; if we previously had an ‘accidental zero’, then we expect $1/\sqrt{2}$. Indeed, this set of measurements provides a reasonably robust answer. This is because $\mathrm{min}(\mathrm{abs}(x-(2y-1)^{2})+\mathrm{abs}(y-(2z-1)^{2}))=1/\sqrt{2}$ when $\alpha=\pi$, providing a significant ‘gap’ to the value expected for $\alpha=0$ (i.e. 00).
@@
.4 IV. Details for Method 2: Consecutive Deflection
Suppose that we start in a state purely in band $\mathinner{|{u_{n}}\rangle}$. When we evolve this in and out of a node at an angle such that the frame has rotated by $\phi^{\mathrm{frame}}_{1}$, upon exiting the node the state is given by
$\cos{\phi^{\mathrm{frame}}_{1}}\mathinner{|{u_{n}}\rangle}-\sin{\phi^{\mathrm{frame}}_{1}}\mathinner{|{u_{n+1}}\rangle}.$ (S27)
@@
$e^{i\alpha^{\prime}}\cos{\phi^{\mathrm{frame}}_{1}}\mathinner{|{u_{n}}\rangle}-\sin{\phi^{\mathrm{frame}}_{1}}\mathinner{|{u_{n+1}}\rangle}.$ (S28)
Upon entering and exiting this second node with a frame rotation angle $\phi^{\mathrm{frame}}_{2}$, we have (denoting $\cos{\phi^{\mathrm{frame}}_{i}}=c_{i},\sin{\phi^{\mathrm{frame}}_{i}}=s_{i}$)
@@
$n_{1}=|e^{i\alpha}c_{1}c_{2}-s_{1}s_{2}|^{2}=\cos^{2}{\left(\phi^{\mathrm{frame}}_{1}+\phi^{\mathrm{frame}}_{2}\right)}+\sin{2\phi^{\mathrm{frame}}_{1}}\sin{2\phi^{\mathrm{frame}}_{2}}\sin^{2}{\left(\alpha/2\right)}.$ (S30)
@@
$n_{1}$ $=\frac{1}{2}\left[\cos{\left(2\phi^{\mathrm{frame}}_{1}+2\phi^{\mathrm{frame}}_{2}\right)}+1+2\sin{2\phi^{\mathrm{frame}}_{1}}\sin{2\phi^{\mathrm{frame}}_{2}}\sin^{2}{\left(\alpha/2\right)}\right]$ (S31)
$=\frac{1}{2}\left[\cos{2\phi^{\mathrm{frame}}_{1}}\cos{2\phi^{\mathrm{frame}}_{2}}-\cos{\alpha}\sin{2\phi^{\mathrm{frame}}_{1}}\sin{2\phi^{\mathrm{frame}}_{2}}+1\right]$ (S32)
$=\frac{1}{2}\left[A\cos{\left(2\phi^{\mathrm{frame}}_{1}-\beta\right)}+1\right]$ (S33)
@@
$A^{2}=\cos^{2}{\left(2\phi^{\mathrm{frame}}_{2}\right)}+\cos^{2}{\left(\alpha\right)}\sin^{2}{\left(2\phi^{\mathrm{frame}}_{2}\right)},$ (S34)
$\tan{\beta}=-\cos{\left(\alpha\right)}\tan(2\phi^{\mathrm{frame}}_{2})$ (S35)
@@
.5 V. Numerical calculations
@@
In this section we present numerical calculations to verify that the phases and band populations are as expected from the theory, in addition to further explanations and examples to further clarify the physics.
We define the path of interest $\bm{k}(t)$, and solve Eqn. S13 numerically using Runge-Kutta [51], for an initial state being a pure band eigenstate at $\bm{k}(0)$. To confirm the results of the interferometry, we must find the phase difference along two paths for a given band. We did this in two different ways. First, if the final states along paths $I$ and $II$ are given by $\mathinner{|{u_{I}}\rangle}$ and $\mathinner{|{u_{II}}\rangle}$, and we would like to find the phase difference of the component in band $\mathinner{|{u_{n}}\rangle}$, we evaluate the argument of $\mathinner{\langle{u_{I}|u_{n}}\rangle}\mathinner{\langle{u_{n}|u_{II}}\rangle}$. For example, $\Delta\alpha_{n+1}-\Delta\alpha_{n}=\mathinner{\langle{u_{I}|u_{n}}\rangle}\mathinner{\langle{u_{n}|u_{II}}\rangle}-\mathinner{\langle{u_{I}|u_{n+1}}\rangle}\mathinner{\langle{u_{n+1}|u_{II}}\rangle}$. This then directly yields the phase differences, which in all cases are as expected. Alternatively, we can follow more closely the experimental procedure. After finding the final states, we construct the wavefunctions for each spin after the $\pi/2$-pulse, and find the corresponding populations. From these, we follow the techniques outlined in Sections III and IV to extract the phase information.
We confirmed that the techniques are robust to experimental errors in the path and accelerations. For method 1, using the technique described in Section II, the gap between expected results for similar and opposite frame charges mean that these can still be reliably distinguished. We generically find that the requirement for adiabaticity away from the nodes provides the strongest constraint, but depends on the precise Hamiltonian. For example, for the Kagome lattice we require $\beta\lesssim 1$. Distinguishing relative charges is consistently reliable for randomly distributed $\delta$ with standard deviation $0.05$ (in units with lattice constant $a=1$), where we have also included random $5\%$ error in accuracy of accelerations. On the other hand, when studying the ‘split gamma’ nodes, we require $\beta\lesssim 0.2$, which accordingly forces smaller errors in $\delta$ of order $0.01$, and due to the longer times the accuracy in acceleration must be within $1\%$. Similar considerations apply to method 2; the primary concern is keeping the dynamic phase evolved between the nodes consistent within each run of the experiment, which also requires maintaining adiabaticity away from the nodes.
The above requirements can be compared to parameters achieved in previous experiments [34]. A difficulty is that decoherence occurs on a timescale of order $\sim 1~{}\mathrm{ms}$, which puts a lower limit on the acceleration $\beta\gtrsim 1$. This indicates that the techniques developed here are right on the limit of what can be achieved with state of the art experiments. The conditions are easiest when dynamic phases cancel along the two paths, since reference interferometry loops need not be performed, making the Kagome lattice an ideal setting for initial experiments. It should also be noted that reference loops can be avoided if the energies along the path are known.
.6 V A. Results and Examples for Method 1
.6.1 Static Kagome with NN and NNN tunneling terms
@@
.6.2 Anomalous Floquet Phases
@@
To deal with the dynamic phases numerically, we employ two approaches. First, we calculate directly the dynamic phase by integrating the band energies along the paths and subtract it at the end to leave only the geometric contribution. We also work directly from the calculated populations, using the reference interferometry technique described in Sec. III “Accounting for Dynamic Phases”, to extract the geometric phase difference. Both methods are numerically confirmed to allow for determining the correct geometric phase difference.
@@
.7 V B: Results and Examples for Method 2
In this section we present corroborating numerical results for Method 2. In particular, we take the opportunity to step-by-step demonstrate how our techniques work when the winding is not uniform, and emphasise the distinction between the frame-rotation angle $\phi^{\mathrm{frame}}$ and the path rotation angle $\phi^{\mathrm{path}}$.
We have confirmed the technique for nearest-neighbour Kagome in Fig. 3 for similarly- and opposite-charged nodes (K and K’). We first note that the value of $\cos{\alpha^{\prime}}$ extracted from the ‘double acceleration’ plots (Fig. 3c) matches up with dynamic phases found by direct integration of energies along the paths, as well as the shifts in the oscillations (Fig. 3b) are quantitatively as expected. In this case, the frames wind uniformly ($\phi^{\mathrm{frame}}=\pm\phi^{\mathrm{frame}}$) and thus choosing $\phi^{\mathrm{frame}}=\pi/2$ fixes $|\phi^{\mathrm{frame}}|=\pi/4$ as required.
@@
Step 1: We first must determine the path angles $\phi^{\mathrm{frame}}$ that lead to frame rotation angles $|\phi^{\mathrm{frame}}|=\pi/4$. To do this, we use the methods of Ref. [34]. Taking the path which joins the nodes as the fixed point, study the population changes upon entering and exiting each node at an angle. This yields the winding shown in Fig. S10. From this data, we can read off the values of $\phi^{\mathrm{frame}}$ for which the population is $0.5$, corresponding to $|\phi^{\mathrm{frame}}|=\pi/4$. In this case, we obtain $\phi^{\mathrm{frame}}_{1}=2.43$ or $\phi^{\mathrm{frame}}_{1}=-0.71$. Repeating this for the other node, we have $\phi^{\mathrm{frame}}_{2}=2.43$ (we do not need the negative option here).
Figure S10: Example of a non-uninorm frame winding around a singularity. The population in band 1 after entering and exiting the node at an angle $\phi^{\mathrm{path}}_{1}$ relative to the path joining the nodes, which corresponds to the frame rotation angle $\phi^{\mathrm{frame}}_{1}$ through $p_{1}=\cos^{2}\phi^{\mathrm{frame}}_{1}$. This allows the value of $\phi^{\mathrm{path}}_{1}$ to be chosen such that the frame rotation angle $|\phi^{\mathrm{frame}}|=\pi/4$. Notice that the winding is not uniform, but the net winding is $\pi$ as it is a topologically protected value.
Step 2: For both these values of $\phi^{\mathrm{frame}}_{1}$, and the fixed $\phi^{\mathrm{frame}}_{2}$, we prepare a wavepacket in band 1 and accelerate it along the required path with an acceleration $2a$ (where $a$ must be chosen such that the evolution is adiabatic), before measuring the final population in band 1. For the two values of $\phi^{\mathrm{frame}}_{1}$ we obtain $p_{1}=0.214$ and $p_{1}=0.786$. The square of the difference of these values gives us the dynamic phase $\cos^{2}{\alpha/2}=0.326$.
Step 3: Finally, we set $\phi^{\mathrm{frame}}_{1}=2.43$ and $\phi^{\mathrm{frame}}_{2}=2.43$ (i.e. the $\phi^{\mathrm{frame}}$’s have the same sign), and accelerate the wavepacket along the path with an acceleration $a$. We find that the final population in band 1 is $0.326$, which is the same as $\cos^{2}{\alpha/2}$, and therefore we know that along this line joining the nodes, the nodes are oppositely charged, and the patch Euler class along this line is 0.